# The accreted stellar haloes of Milky Way-mass galaxies as a probe of the nature of the dark matter

### v1

@@ 1 Introduction
Structure formation within the $\Lambda$CDM model proceeds in a hierarchical fashion (Davis et al., 1985), whereby small structures form first and then grow larger through mass accretion and mergers. This process results in the formation of virialised structures, known as dark matter haloes, that follow quasi-universal density profiles (Navarro et al., 1997) over twenty orders of magnitude in halo mass (Wang et al., 2020). Haloes can grow in mass discretely by the accretion of neighbouring dark matter haloes, which are not necessarily disrupted. Their subsequent evolution after being accreted depends on their mass relative to their host, their orbital and structural parameters, and the nature of the dark matter itself (e.g. halos could evaporate as they merge if the dark matter has a non-zero interaction cross-section; Kummer et al. 2018). Those sufficiently massive to experience dynamical friction (Chandrasekhar, 1943) lose orbital angular momentum and energy, eventually disrupting in the core of the host. The remaining objects are primarily affected by tidal stripping and shock heating. The loss of mass that occurs during merging and stripping can lead to the formation of streams and shells of material, leaving an imprint in phase space that can remain long after they first formed.
@@
Since the formation of the stellar halo is intimately related to the stripping and disruption of satellite galaxies, changing how many are accreted and how efficiently they are stripped from stars may result in noticeable differences in the properties of stellar haloes. For example, the suppression of the least massive satellites could affect the number of accreted stars in the outer stellar halo, as these are the progenitors that dominate its outskirts (e.g. Fattahi et al., 2020). Additionally, changing the structural parameters of haloes hosting dwarf galaxies can alter the spatial distribution and kinematics of the resulting remnant (Amorisco, 2017; Vasiliev et al., 2022).
Given the potential importance of the assumed nature of DM on the properties of stellar halos, it is surprising that few studies have explored the connection between the two in detail. Beyond stream ‘gapology’ (e.g. Ibata et al., 2002) and stellar wakes (e.g. Buschmann et al., 2018), the most recent example focused on the stellar haloes around dwarf galaxies using idealised N-body simulations (Deason et al., 2022). However, this mass scale is subject to substantial theoretical (e.g. Brook et al., 2014; Read et al., 2017; Jethwa et al., 2018; Graus et al., 2019; Benitez-Llambay & Frenk, 2020) and observational uncertainties (e.g. Kazantzidis et al., 2011; Sestito et al., 2023; Waller et al., 2023).
@@ 2 Simulations
In this section, we describe the simulation code, galaxy formation physics and alternative dark matter models used in our simulations.
@@ 2.1 The galaxy formation model
2.2 Warm dark matter
We obtain the power spectrum of WDM, $P_{\mathrm{WDM}}(k)=T^{2}(k)P_{\mathrm{CDM}}(k)$, using the transfer function of Bode et al. (2001):
$T^{2}(k)=[1+(\alpha k)^{2\nu}]^{-5/\nu}\,.$ (1)
Here, $\nu$ is a fitting constant equal to 1.2 and the parameter $\alpha$ depends on the assumed mass of the WDM particle:
$\alpha=0.049\Big{[}\dfrac{m_{\mathrm{th}}}{\mathrm{keV}}\Big{]}^{-1.11}\Big{[}\dfrac{\Omega_{\mathrm{WDM}}}{0.25}\Big{]}^{0.11}\Big{[}\dfrac{h}{0.7}\Big{]}^{1.22}\,h^{-1}\,\mathrm{Mpc}\,.$ (2)
For this work we assume $m_{\mathrm{th}}=2.5\,\mathrm{keV}$. This is lighter than the equivalent thermal relic mass of a 7 keV sterile neutrino model attributed through its decay to the unidentified 3.5 keV X-ray line (Boyarsky et al., 2014). Nonetheless, we chose this value to enhance the differences with respect to CDM to allow for an easier comparison. As a reference point, the Jean’s mass of a perturbation with a wavelength equal to the one where the WDM power spectrum is half of the CDM one is $m_{1/2}=1.4\times 10^{9}\,\mathrm{M}_{\odot}$. This corresponds to the mass scale where the differences with respect to CDM are expected to be noticeable.
2.3 Self-interacting dark matter
Self-interactions are modelled using the Monte-Carlo implementation described in Robertson et al. (2017). Dark matter particles can scatter with each other when they are closer than the gravitational softening length of the simulations, $\epsilon$. The probability of any two neighbouring particles scattering, within a timestep of size $\Delta t$, is a function of their relative velocity and the assumed cross-section, $\sigma_{\mathrm{SIDM}}$:
$P_{ij}=\dfrac{\sigma_{\mathrm{SIDM}}|\vec{v}_{i}-\vec{v}_{j}|}{\dfrac{4\pi\epsilon^{3}}{3}}$ (3)
In this study, we use a constant isotropic cross-section of $\sigma_{\mathrm{SIDM}}=10~{}\mathrm{cm}^{2}\mathrm{g}^{-1}$.
@@ 3 Methods
Here we discuss how we construct the catalogue of dark matter haloes and galaxies in our simulations. We also explain how we link them across time to build their merger trees, and how this information is used to identify the accreted stellar halo and its progenitor galaxies. We also discuss how we remove the spurious structures that form in WDM (e.g. Wang & White, 2007), and how we select our sample of host galaxies.
@@ 3.2 Spurious group removal
Particle-based simulations with a resolved power spectrum cut-off result in the spurious fragmentation of filaments due to the discrete representation of the underlying density field (Wang & White, 2007). This artificially boosts the number of identified structures below the mass scale where structure formation is strongly suppressed. We therefore clean the WDM catalogues from these structures using the criteria of Lovell et al. (2014).
Firstly, we remove all subhaloes whose peak bound mass is lower than the scale where spurious structures are expected to dominate:
$M_{\rm lim}=5.05\,\bar{\rho}dk^{-2}_{\rm peak},$ (4)
where $d$ is the mean interparticle separation, $k_{\rm peak}$ the wavelength at which the dimensionless power spectrum peaks, and $\bar{\rho}$ is the mean density of the universe. For our setup, this corresponds to $M_{\rm lim}=1.4\times 10^{8}\,\,\mathrm{M}_{\odot}$. Lastly, we compute the sphericity ($s\equiv c/a$, where $c$ and $a$ are the smallest and largest eigenvalues of the inertia tensor) of the Lagrangian region associated to the proto-haloes of the remaining subhaloes, using the particles bound to them when they first reached their half peak mass. Those with significantly flattened distributions ($s\leq 0.16$), and hence likely spurious, are removed.
@@ 4.1 Stellar halo masses
We begin by analysing how the accreted stellar mass varies across the galaxies in our sample, and examine if there is any dependence on the assumed model of dark matter. The ex-situ stellar masses for the galaxies formed in the CDM simulation are shown along the horizontal axis of Fig. 1, with the median and 16th - 84th percentiles of our sample being $M^{\mathrm{CDM}}_{\mathrm{exsitu}}=2.1^{+2.3}_{-0.7}\times 10^{9}\,\,\mathrm{M}_{\odot}$. Although roughly consistent with observational estimates for the mass of the MW stellar halo, we note that our quoted value does not include the in-situ stellar halo. Hence, the masses we measure here represent a lower limit of total mass of our simulated stellar haloes.
@@
We further note that half of our sample has an accreted stellar halo more massive than the total mass of the Milky Way stellar halo. As our simulated Milky Way analogues are only selected based on their present-day virial mass, it is likely that the galaxies in our sample formed later than counterparts whose assembly histories more closely match that of our Galaxy. As the mass of stellar haloes correlate with their time of formation (e.g. Deason et al., 2019), one might thus expect more massive stellar haloes in our sample. We further note that ex-situ mass fraction in our sample, defined here as the ratio of the ex-situ mass to the total stellar mass within a 30 kpc spherical aperture, is $\sim 21\%$ and hence significantly larger than the inferred $\sim 1\%$ for our Galaxy. Aside from the tendency for the accreted mass to be somewhat higher, this is also driven by the fact the EAGLE model of galaxy formation underpredicts the stellar masses of central galaxies that are hosted in Milky Way mass haloes.
@@
Shifting our focus to the impact of the assumed nature of dark matter on the properties of the accreted stellar haloes, Fig 1 indicates that there are up to $\sim$20%-level differences at a fixed stellar halo between the different DM models. Despite a tendency toward lower accreted masses, with a median ratio and scatter of $M^{\mathrm{WDM}}_{\mathrm{exsitu}}/M^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.9^{+0.2}_{-0.2}$ and $M^{\mathrm{SIDM}}_{\mathrm{exsitu}}/M^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.9^{+0.1}_{-0.1}$, the population is consistent with those found in CDM. The relatively small differences across the various DM models explored here indicates that the nature of DM plays a minor role in setting the overall ex-situ mass of the stellar halo. This is because the largest progenitors, which contribute most of the accreted mass, have similar stellar masses across the models, as shown in the appendix (Fig. 7). Nonetheless, it is worth remembering that we have not considered the in-situ component of the stellar halo. As such, our mass estimate corresponds to a lower bound of the total stellar halo mass, and does not imply that the total (ex-situ plus in-situ) remains similar across DM models.
@@
We conclude that the total accreted mass of stellar haloes is largely insensitive to changes caused by the nature of the dark matter. The reason behind this similarity is the fact that neither warm nor self-interacting dark matter change the mass or abundance of the largest stellar halo progenitors. Since these dominate the mass budget of the stellar halo, their ex-situ masses do not change substantially across models. However, this does not preclude the presence of differences on a spatial or dynamical level, which we explore next.
@@ 4.2 Stellar density profiles
We explore this in the top panel of Fig. 2, which shows the median density profiles obtained from our sample of MW-mass haloes. We measure the profiles for each individual galaxy using 30 spherical shells whose edges are logarithmically spaced between $0.005R_{200}$ and $R_{200}$. We define the origin as the centre of mass of all dark matter particles within 5 kpc from the centre of potential of the main galaxy, as identified by SUBFIND. The spatial offset between this centre and the halo centre of potential (or the centre of mass of the stars) is comparable to the gravitational softening of the simulation.
Figure 2: Top panel: median 3D density profiles of the ex-situ (dashed) and in-situ (dotted) stellar components around MW-mass haloes, as well as its DM component (solid). This is shown across their CDM (black), WDM (red) and SIDM (blue) variants. Middle panels: median ratio of the ex-situ stellar and DM density at a given spherical shell across DM variants, relative to their values in the CDM counterparts. Bottom panel: median differential ex-situ density relative to the total stellar density, as a function of distance to the centre. The distributions shown in the bottom three panels have been smoothed using a linear Savitzky–Golay filter over three consecutive bins, and the shaded regions indicate the 16th to 84th percentiles.
@@
Unsurprisingly, the in-situ stellar and dark matter density profiles of the WDM model are consistent with the CDM counterparts across the whole radial range we consider. We show these density profiles in the top panel of Fig. 2, which we measured using the same binning scheme as that employed for the accreted component. The similarity in the density distribution of the central galaxy and its host dark matter halo is expected. The mass scale affected by our chosen cut-off in the power spectrum is orders of magnitude lower than MW-mass scales. This means that the formation of the host DM halo is not affected by a delay in its formation, and so its structural parameters are largely the same in CDM and WDM. As the galaxy formation model is the same, the galaxies that form at its centre are very similar.
@@
Shifting to the SIDM model, it is apparent that the outskirts of the ex-situ density profiles are similar to those in CDM. However, contrary to what we find in WDM, the ex-situ stellar density is strongly suppressed in the central $0.05R_{200}$. The radial dependence of the accreted density is also different than the corresponding CDM and WDM counterparts, appearing significantly flatter towards the centre. To quantify how much the slope has changed, we fit a power law, $\rho\propto r^{\alpha}$, to the ex-situ profiles within $0.02R_{200}$. We do this for both CDM and SIDM, measuring a median value in our sample of $\langle\alpha_{\mathrm{CDM}}\rangle=-1.7^{+0.3}_{-0.4}$ and $\langle\alpha_{\mathrm{SIDM}}\rangle=-1.4^{+0.4}_{-0.3}$. Despite the large scatter in the slopes, the accreted stellar mass profiles are systematically flatter in SIDM than in CDM. Beyond the slope of the profiles, the central ($r\sim 0.01R_{200}$) ex-situ density decreases to a value of $\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.3^{+0.9}_{-0.2}$ relative to the CDM counterparts, with some stellar haloes having an order of magnitude lower in density.
@@
As stellar haloes are tracers of the underlying gravitational potential field, we explore how the changes in the spatial distribution of accreted stars in SIDM compare to the DM halo of their host galaxy. The median ratio of the DM density profiles in SIDM relative to CDM is plotted in the second-to-last panel of Fig. 2. Here we see that the suppression in the DM density at $0.01R_{200}$ ($\rho^{\mathrm{SIDM}}_{\mathrm{DM}}/\rho^{\mathrm{CDM}}_{\mathrm{DM}}=0.2^{+0.2}_{-0.1}$) is stronger and less variable than that observed in the accreted stellar densities. The scale on which the expansion is present in both the DM and accreted stellar components is similar.
@@
We also highlight the fact that the in-situ stellar component exhibits differences relative to their counterparts in CDM and WDM (see top panel of Fig. 2). In particular, their central densities are lower than the corresponding galaxies formed in CDM and WDM, with a slight enhancement in the outskirts of the galaxy ($r\sim 0.01$ to $0.1R_{200}$). As this is likely partly driven by the shallower potential well present in the host, it will be interesting to explore in the future how the perturbative effect of mergers and fly-bys differ in the formation process of in-situ stellar haloes in SIDM cosmologies.
@@
Nonetheless, the magnitude and shape of the ex-situ and in-situ expansion differ between each other, leading to different local ex-situ mass fractions, $f_{\mathrm{exsitu}}=\rho_{\mathrm{exsitu}}/\rho_{*}$, as a function of radius. The median values measured across our sample are shown in the bottom panel of Fig. 2, where we see that the radial dependence in SIDM is substantially different to CDM and WDM. Its value is always lower in the spatial range where the halo has expanded ($r<0.05R_{200}$), and retains a constant value within $\sim 0.02R_{200}$ instead of increasing like in CDM and WDM.
@@
We confirm that the way in which accreted stellar material is distributed within $0.05R_{200}$ varies depending on the assumed dark matter model. For WDM (SIDM), the inner regions of the accreted stellar halo are more (less) dense than in CDM, with the most extreme examples having $\rho^{\mathrm{WDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=2.56$ and $\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.05$ at $0.01R_{200}$. Together with the fact that the overall ex-situ mass remains similar across models, these changes suggest a rearrangement of the accreted material, and hence how spatially concentrated it is distributed. Since the spatial distribution of stars reflects their underlying orbit distribution, these differences motivate a closer look at their velocity distributions. Additionally, phase-space signatures originating from major accretion and stripping events might differ relative to CDM, reflecting the change in the rate with which bound mass is lost and the number of accretions that result from the alternative DM variants used in this work. We explore this in the following subsection.

### v2

@@ 1 Introduction
[…]
@@
Since the formation of the stellar halo is intimately related to the stripping and disruption of satellite galaxies, changing how many are accreted and how efficiently they are stripped of stars may result in noticeable differences in the properties of stellar haloes. For example, the suppression of the least massive satellites could affect the properties of the outer stellar halo, either in terms of its extent, mass or the amount of tidal-induced features, as these are the progenitors that dominate its outskirts (e.g. Fattahi et al., 2020). Additionally, changing the structural parameters of haloes hosting dwarf galaxies can alter the spatial distribution and kinematics of the resulting remnant (Amorisco, 2017; Vasiliev et al., 2022).
Given the potential importance of the assumed nature of DM on the properties of stellar halos, it is surprising that few studies have explored the connection between the two in detail. Beyond stream ‘gapology’ (e.g. Ibata et al., 2002) and stellar wakes (e.g. Buschmann et al., 2018), only a limited number of recent studies have explored what effect changing the DM model has on stellar haloes. The most recent example focused on the stellar haloes around dwarf galaxies using idealised N-body simulations (Deason et al., 2022). However, this mass scale is subject to substantial theoretical (e.g. Brook et al., 2014; Read et al., 2017; Jethwa et al., 2018; Graus et al., 2019; Benitez-Llambay & Frenk, 2020) and observational uncertainties (e.g. Kazantzidis et al., 2011; Sestito et al., 2023; Waller et al., 2023). On Milky Way-mass scales, Dooley et al. (2016) used particle tagging on N-body simulations to predict how the outer ($\geq 30~\mathrm{kpc}$) stellar halo changed between CDM and SIDM, finding virtually no differences. However, an important caveat in their study is that they only focused on the outer stellar halo, which is thought to be dominated by low mass progenitors of the stellar halo. Thus, the possibility of differences in the inner stellar halo remain, so long as the most massive progenitors (those whose debris dominate the inner regions) are affected by changes in the DM.
@@ 2 Simulations
In this section, we briefly describe the simulation code, galaxy formation physics and alternative dark matter models used in our simulations. Further details are available in Forouhar Moreno et al. (2022).
@@ 2.1 The galaxy formation model
2.2 Alternative dark matter models
We run two simulations in which the dark matter model differs from CDM, but whose subgrid physics and phases in the initial conditions remain the same. Any changes in the properties of stellar haloes of our simulated galaxy sample are thus primarily caused by differences driven by the nature of dark matter.
The first variation we consider is a warm dark matter model, in which we truncate the initial power spectrum of density fluctuations using the transfer function of Bode et al. (2001). We assume a $2.5\,\mathrm{keV}$ thermal relic particle, which corresponds to a mass scale where differences start to appear relative to CDM (its half-mode mass) of $m_{1/2}=1.4\times 10^{9}\,\mathrm{M}_{\odot}$. Although this mass is lighter than current observational constraints, we choose this mass to enhance differences relative to a CDM model.
The second variation in the dark matter model is through the addition of self-interactions between DM particles, modelled using the Monte-Carlo implementation of Robertson et al. (2017). The cross-section we use is velocity-independent and equals $\sigma_{\mathrm{SIDM}}=10~\mathrm{cm}^{2}\mathrm{g}^{-1}$. We use the same initial conditions as those generated for the CDM simulation.
@@ 3 Methods
Here we summarise the methods used in Forouhar Moreno et al. (2022) to construct the catalogue of dark matter haloes and galaxies, as well as how they are linked across time, and how spurious subhaloes are removed from the WDM simulation. We also explain how merger trees are used to identify the accreted stellar halo and its progenitor galaxies.
@@ 3.2 Spurious group removal
Particle-based simulations with a resolved power spectrum cut-off result in the spurious fragmentation of filaments due to the discrete representation of the underlying density field (Wang & White, 2007). To remove the resulting spurious subhaloes, we use the method of Lovell et al. (2014). First, we remove all subhaloes whose mass is lower than the mass scale where spurious structures are expected to dominate ($M_{\rm lim}=1.4\times 10^{8}\,\,\mathrm{M}_{\odot}$ for our setup). We then remove all remaining subhaloes whose proto-halo Lagrangian region is highly flattened, i.e. with a sphericity $s$ lower than 0.16 ($s\equiv c/a$, where $c$ and $a$ are the smallest and largest eigenvalues of the inertia tensor).
@@ 4.1 Stellar halo masses
We begin by analysing how the accreted stellar mass varies across the galaxies in our sample, and examine if there is any dependence on the assumed model of dark matter. The ex-situ stellar masses [^3] Measured as the total stellar mass of ex-situ star particles that are enclosed by a spherical aperture of 100 kpc placed on the centre of the galaxy. for our sample of galaxies in the CDM simulation are shown along the horizontal axis of Fig. 1, with the median and 16th - 84th percentiles of our sample being $M^{\mathrm{CDM}}_{\mathrm{exsitu}}=2.0^{+1.8}_{-0.7}\times 10^{9}\,\,\mathrm{M}_{\odot}$. Although roughly consistent with observational estimates for the mass of the MW stellar halo, we note that our quoted value does not include the in-situ stellar halo. Hence, the masses we measure here represent a lower limit of total mass of our simulated stellar haloes.
@@
We further note that half of our sample has an accreted stellar halo more massive than the total mass of the Milky Way stellar halo. As our simulated Milky Way analogues are only selected based on their present-day virial mass, it is likely that the stellar halo around the galaxies in our sample formed later than counterparts whose assembly histories more closely match that of our Galaxy. There are two reasons that explain this expectation. First, galaxies hosting GE/S-like features in their stellar haloes form earlier than galaxies that do not exhibit such a radially-anisotropic component (Fattahi et al., 2019). Hence, our mass selection will contain a higher number of ‘average’ MW-mass haloes without a highly radially anisotropic component. Second, the stellar halo of the real Milky Way formed earlier than a mass-selected sample of simulated MW-mass galaxies (e.g. Deason et al., 2019), which also means that we might expect more massive stellar haloes (relative to the stellar mass of their host galaxy) in our sample. Indeed, the ex-situ mass fraction in our sample, defined here as the ratio of the ex-situ mass to the total stellar mass within a 30 kpc spherical aperture, is $\approx 21\%$ and hence significantly larger than the inferred $\approx 1\%$ for our Galaxy.
@@
Shifting our focus to the impact of the assumed nature of dark matter on the properties of the accreted stellar haloes, Fig 1 indicates that there are up to $\approx$20%-level differences at a fixed stellar halo mass between the different DM models. Despite a tendency toward lower accreted masses, with a median ratio and scatter of $M^{\mathrm{WDM}}_{\mathrm{exsitu}}/M^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.9^{+0.2}_{-0.2}$ and $M^{\mathrm{SIDM}}_{\mathrm{exsitu}}/M^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.8^{+0.1}_{-0.1}$, the populations are roughly consistent with those found in CDM. The relatively small differences across the various DM models explored here indicate that the nature of DM plays a minor role in setting the overall ex-situ mass of the stellar halo. This is because the largest progenitors, which contribute most of the accreted mass, have similar stellar masses across the models, as shown in the appendix (Fig. 7). Nonetheless, it is worth remembering that we are not including the in-situ stars that have halo-like orbits in our analysis (i.e. the in-situ stellar halo). As such, our mass estimate corresponds to a lower bound of the total stellar halo mass. A future avenue to explore is whether or not larger amounts of in-situ stars have halo-like orbits in alternative dark matter models. One might expect the fraction to change owing to the shallower gravitational potentials in SIDM, making their central galaxies more susceptible to tidal perturbations.
@@
We conclude that the total accreted mass of stellar haloes is largely insensitive to changes caused by the nature of the dark matter. The reason behind this similarity is that, as we will show explicitly in §4.4, neither warm nor self-interacting dark matter change the mass or abundance of the most massive stellar halo progenitors. Since these dominate the mass budget of the stellar halo, their ex-situ masses do not change substantially across models. This is perhaps not a surprising fact, as it is known that the stellar masses of the surviving population of satellite galaxies do not change significantly across DM models. Nonetheless, up until this study there had been no explicit verification of whether the stellar masses of disrupted satellites also remain relatively unchanged across different DM models. For example, one could have expected that differences in the formation time of WDM subhaloes could have reduced their stellar mass (e.g. if the time subhaloes have from when they are massive enough to form stars until they are accreted and quenched becomes significantly shorter). We have shown that, at least for the models and mass ranges we consider here, this is not the case.
That being said, the stellar masses of the low mass end of disrupted satellites may be more sensitive to the suppression and delays in the formation of structure. If so, smaller scale structure that is part of the stellar halo may retain an imprint of the power-spectrum cut-off, e.g. in the number and properties of stellar streams. Similarly, velocity-dependent SIDM models that predict gravothermal collapse in the lowest mass subhaloes could also affect small scale stellar halo properties. Leaving aside the stellar masses, the lack of differences does not preclude changes on a spatial or dynamical level, which we explore next.
@@ 4.2 Stellar density profiles
We explore this in the top panel of Fig. 2, which shows the median density profiles obtained from our sample of eight MW-mass haloes. We measure the profiles for each individual galaxy using 30 spherical shells whose edges are logarithmically spaced between $0.005R_{200}$ and $R_{200}$. We define the origin as the centre of mass of all dark matter particles within 5 kpc from the centre of potential of the main galaxy, as identified by SUBFIND. The spatial offset between this centre and the halo centre of potential (or the centre of mass of the stars) is comparable to the gravitational softening of the simulation.
Figure 2: Top panel: median of the 3D density profiles of the ex-situ (dashed) and in-situ (dotted) stellar components around the eight MW-mass haloes in our sample, as well as its DM component (solid). Note that the in-situ stellar density also includes the stellar disc of the galaxy. This is shown across their CDM (black), WDM (red) and SIDM (blue) variants. Middle panels: median ratio of the ex-situ stellar and DM density at a given spherical shell across DM variants, relative to their values in the CDM counterparts. Bottom panel: median differential ex-situ density relative to the total stellar density, as a function of distance to the centre. The distributions shown in the bottom three panels have been smoothed using a linear Savitzky–Golay filter over three consecutive bins, and the shaded regions indicate the 16th to 84th percentiles.
@@
Unsurprisingly, the median in-situ stellar and dark matter density profiles of the WDM model are consistent with the CDM counterparts across the whole radial range we consider. We show these density profiles in the top panel of Fig. 2, which we measure using the same binning scheme as that employed for the accreted component. The similarity in the density distribution of the central galaxy and its host dark matter halo is expected. The mass scale affected by our chosen cut-off in the power spectrum is orders of magnitude lower than MW-mass scales. This means that the formation of the host DM halo is not affected by a delay in its formation, and so its structural parameters are largely the same in CDM and WDM. As the galaxy formation model is the same, the galaxies that form at its centre are very similar.
@@
Shifting to the SIDM model, it is apparent that the outskirts of the ex-situ density profiles are similar to those in CDM. The lack of differences in the outer stellar halo between SIDM and CDM was also highlighted by the N-body simulations used in Dooley et al. (2016). However, their study only explored the stellar halo density profiles beyond 30 kpc, which corresponds to approximately $(0.1-0.15)\times R_{\mathrm{200}}$, depending on the assumed value of $R_{200}$. Indeed, we find that the ex-situ stellar density is strongly suppressed in the central $0.05R_{200}$ in the SIDM model relative to the CDM one. The radial dependence of the accreted density in SIDM is also different to the corresponding CDM and WDM counterparts, appearing significantly flatter towards the centre. To quantify how much the slope has changed, we fit a power law, $\rho\propto r^{\alpha}$, to the ex-situ profiles within $0.02R_{200}$ for each of the eight MW-mass galaxies in our sample. We do the fitting for the stellar haloes in the CDM and SIDM models, measuring a median value of $\langle\alpha_{\mathrm{CDM}}\rangle=-1.7^{+0.3}_{-0.4}$ and $\langle\alpha_{\mathrm{SIDM}}\rangle=-1.4^{+0.4}_{-0.3}$. Despite the large scatter in the slopes, the median accreted stellar mass profiles are systematically flatter in the central regions for stellar haloes in the SIDM model relative to CDM. Beyond the slope of the profiles, the central ($r\approx 0.01R_{200}$) ex-situ density decreases to a value of $\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.3^{+0.9}_{-0.2}$ relative to the CDM counterparts, with some stellar haloes having an order of magnitude lower density.
@@
As stellar haloes are tracers of the underlying gravitational potential field, we explore how the changes in the spatial distribution of accreted stars in SIDM compare to the DM halo of their host galaxy. The median ratio of the DM density profiles in SIDM relative to CDM is plotted in the second-to-last panel of Fig. 2. Here we see that the suppression in the DM density at $0.01R_{200}$ ($\rho^{\mathrm{SIDM}}_{\mathrm{DM}}/\rho^{\mathrm{CDM}}_{\mathrm{DM}}=0.2^{+0.2}_{-0.1}$) is stronger and has considerably less scatter than the suppression we measure in the accreted stellar densities. The spatial scale on which the expansion is present in both the DM and accreted stellar components is similar.
@@
We also highlight the fact that the in-situ stellar component (which also includes the stellar disc) exhibits differences relative to their counterparts in CDM and WDM (see top panel of Fig. 2). In particular, their central densities are lower than the corresponding galaxies formed in CDM and WDM, with a slight enhancement in the outskirts of the galaxy ($r\approx 0.01$ to $0.1R_{200}$). As this is likely partly driven by the shallower potential well present in the host, it will be interesting to explore in the future how the perturbative effect of mergers and fly-bys differ in the formation process of in-situ stellar haloes in SIDM cosmologies.
@@
Despite the fact that the ex-situ and in-situ stellar densities are suppressed in SIDM relative to CDM, the magnitude and shape of the suppression differ. Consequently, the local ex-situ mass fraction, $f_{\mathrm{exsitu}}=\rho_{\mathrm{exsitu}}/\rho_{*}$, differs in SIDM from CDM . The median values measured across our sample of eight MW-mass galaxies are shown in the bottom panel of Fig. 2, where we see that the radial dependence in SIDM is substantially different to CDM and WDM. Its value is always lower in the spatial range where the halo has decreased its density compared to CDM ($r<0.05R_{200}$), and retains a constant value within $\approx 0.02R_{200}$ instead of increasing with radius like in CDM and WDM.
@@
We confirm that the way in which accreted stellar material is distributed within $0.05R_{200}$ varies depending on the assumed dark matter model. For WDM (SIDM), the inner regions of the accreted stellar halo are more (less) dense than in CDM, with the most extreme examples having $\rho^{\mathrm{WDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=2.56$ and $\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}=0.05$ at $0.01R_{200}$. Together with the fact that the overall ex-situ mass remains similar across models, these changes suggest a rearrangement of the accreted material, and hence how spatially concentrated it is distributed. Since the spatial distribution of stars reflects their underlying orbital distribution, these differences motivate a closer look at their velocity distributions. Additionally, phase-space signatures originating from major accretion and stripping events might differ relative to CDM, reflecting the change in the rate with which bound mass is lost and the number of accretions that result from the alternative DM variants used in this work. We explore this in the following subsection.